UARS Weekly Status Report

13 July 2001




********************************************************************

                UARS WEEKLY STATUS REPORT - 13 JUL 2001

********************************************************************

***********
SOLSTICE
***********

Nothing received.

***********
SUSIM
***********

SUSIM has completed its daily solar scans this past week during periods
of predicted UARS telemetry.

***********
ACRIM
***********

Nothing received.

***********
PEM
***********

***** ALL PEM COMMUNICATION SHOULD NOW BE THROUGH THE PEM2K:: NODE *****

     The PEM team reports that the PEM:: node computer has failed and has
been taken off line.  PEM:: appears to have hardware problems that take the 
system down every few hours.  Automatic transfer NRT tests to PEM2K from
both CDHF2 and CDHFB have succeed and the NRT files have been verified against 
those captured by PEM, which confirm the NRT process is correct.  Thus, PEM 
confirms NRT data from CDHFB.  Thanks to Steve Bove for his work assisting 
PEM with this process.

     PEM/AXIS was activated in its half power mode on June 24 and deactivated
on June 27.  The PEM/AXIS was reactivated in full power mode on July 3.  At 
the current time, AXIS is not generating energy deposition.  SwRI is looking 
into this.  Tests of AXIS production code on the rehost computer are continuing.

     At the present time, PEM energy deposition from its spectrometers is not 
being generated.  Recent data configuration changes have affected PEM software 
causing multiple failures.  These failures are being reviewed by the PEM team.
In order to correct for new telemetry requirements, PEM will need to update 
processing code with a patch delivery.  The delivery patch may include changes 
which would allow the PEM software to execute on a DEC Alpha machine.  GSFC is
helping PEM with some DEC Alpha particulars which have caused PEM level 3
code to fail on DEC Alpha machines (software below level 3 executes on the
DEC Alpha).

    Updated HEPS calibration parameters have been finalized and are ready for 
delivery.  Updated MEPS quality calibration files have been finalized and are 
ready for delivery.  Level 3 software errors, discovered by Dr. Dean Pesnell, 
have been fixed; his suggestions have been incorporated with in the PEM 
production code.  Both updated HEPS and MEPS calibration data, along with the 
Level 3 code changes, will influence the in-situ energy deposition predictions,
so new versions of PEM energy deposition Level 3 files will be generated.

*Beginning of power orbit day (PEM activation): array current >5 amps, array 
     voltage >53.5 volts, and battery current >2.0 amps.  
 End of power orbit day (PEM deactivation): time in power orbit day >4 minutes,
     array current <9 amps, and battery currents <0.2 amps.
 Power orbit day variation: ~39-40 min at yaw to ~50-51 min at maximum beta.

***********
HALOE 
***********

HALOE continues to operate successfully.  All engineering parameters are 
within nominal operating range.  HALOE was powered off at the 
beginning-of-day today as scheduled.  Special calibration sequences were 
commanded during some sunrise and sunset events to verify instrument 
alignment and end to end gain.

***********
MLS
***********

Nothing received.

***********
HRDI
***********

Nothing received.

***********
WINDII
***********

Nothing received.

***********
MPG
***********
 
Daily processing of instrument Daily Activity Plans 
continues with no problems.

***********
CDHF
***********

DAAC DATA:
 - CLAES (output 9) created in August, 1999 has been loaded to the DAAC 
      Transfer Tables.
 - UKMO,NMC and HALOE created in May, 1999 has been previewed in
      preparation for sending to the DAAC.

UARS CDHF REHOST:
 -  Current day production processing continues to be processed only on the
    rehost system and is going well.
 -  NRT was moved July 12 to CDHFB.
                  
***********
FOT
***********

                           REPORTING PERIOD

This report covers 30 June 2001 (Orbit 53581, GMT Day 181) through 06 July 2001 
(Orbit 53685, GMT Day 187).

                         SPACECRAFT OPERATION

The observatory is now in Forward Flight (Southern Hemisphere viewing) and is 
performing nominally under two-battery operations.  The instrument operational 
changes were:


                     INSTRUMENT OPERATIONAL CHANGES


         Instrument     Time       Orbit     Comment
         SOLSTICE     182/1138     53603     Power On
         SUSIM        182/1415     53604     Power On
         ACRIM        182/2116     53609     Power On
         PEM          183/2339     53625     AXIS On
         HALOE        184/2257     53640     Power On


The beta angle ranged between the angles of 5.6 degrees to 29.3 degrees for 
this report period.  The beta angle is now increasing towards a maximum angle 
of 78 degrees on 20 July 2001 (DOY 201).

Spacecraft battery 1, 2, and 3 performance monitoring continues.  Battery 1 
remains OFF the charge relay.  Battery 2 and Battery 3 are currently 
maintaining greater than 23.8 V end-of-night (EON) load bus voltage.  Battery 
temperatures are stable with a temperature delta between Battery 2 and Battery 
3 of 1.2 to 1.6 degrees C.  Additional battery performance data for each 
battery on 29 June 2001 and 6 July 2001 are provided in Appendix D.

The clock error ranged from 8.9 to 0.6 msec during this report period.
Clock rate adjustments are listed in Appendix C.

All instrument and subsystem engineers were advised on the initial excursion of 
ALERTS or Out-Of-Limits conditions detected by the FOT for this report.  A 
summary is listed below and the details of ALERT and Out-Of-Limits occurrences 
are listed in Appendix B.

     OUT-OF-LIMITS  =  00
     ALERTS         =  00

The following real-time command sheets were executed for the listed operational 
element(s) during this report period (see Appendix A).

     FOT            =  09
     HALOE          =  03
     SOLSTICE       =  01
     ACRIM          =  01
     SUSIM          =  04
     HRDI           =  02
     PEM            =  02

UARS Anomaly Reports initiated during this report period are listed in Appendix 
E and summarized below.

     HALOE          =  01
     FOT            =  01


One (1) Anomaly Report remained OPEN at the end of this report period (see 
Appendix E).

TRANSPONDER FREQUENCIES:
 Transponder A:  00/316  - 2287.496493
 Transponder B:  01/119  - 2287.499116


                     GROUND SYSTEM OPERATION

Two (2) operational support problems occurred this period.  Anomaly Reports are 
included in Appendix E.

The data loss calculations supplied by Data Capture Facility (DCF) from mission 
start through UTC Day 01/186 show the observatory data loss to be 302 hours, 44 
minutes, 57.624 seconds. This shows a loss of 1 hour, 04 minutes and 12.784 
seconds since the last report period. This is a 0.39 percent data loss, which 
equals a 99.61 percent data capture for the mission.

The increase of 1 hour, 04 minutes and 12.784 seconds occurred over five (5) 
days:
 1 hour and 40.592 seconds on Day 180 due to Station preserving power for 
YAW maneuver.
 --No science data was lost
 2 minutes and 07.488 seconds on Day 181 due to No Data -- Overlap
 37.632 seconds on Day 182 due to a Data Hit
 47.072 seconds on Day 183 due to a Data Hit


                      OTHER SIGNIFICANT EVENTS

                      UARS SATELLITE OPERATIONS

An instrument anomaly was experienced on 05 May 2001.  The instrument involved 
was the ACRIM II instrument.  All ACRIM telemetry for currents and voltages all 
dropped to zero values suddenly.  The ACRIM team in coordination with the 
UARSFOT conducted analysis and command tests to determine the cause of the 
anomaly.  Analysis shows that the anomaly was caused by failed circuitry in the 
ACRIM electronics box.  The FOT is awaiting a future plan of action from the 
ACRIM team.

The software test team, in conjunction with the FOT, has completed a large 
portion of the testing for the new front end processor.  The new TAC still has 
yet to undergo DSN and realtime TDRS testing, but has completed simulations of 
receipt of telemetry, commanding, and NCC/GCMR testing.  DSN testing is 
scheduled for 18 July 2001.

UARSFOT will also begin supporting testing of the rehosted UARS ground system.  
A prototype system has been built in accordance with the UARS documentation and 
operational database.  A meeting will be held on 10 July to determine the test 
plan.

UARS attitude has been adjusted over the course of the last few yaw cycles by 
loading new alignment matrices for the TAM.  Good alignments have been found 
for both directions of flight and the performance of the TAM should remain 
stable and accurate. 


                              MEETINGS

The FOT held the UARS Three TDRS Kickoff Meeting on 03 July 2001.  This meeting 
was established to track the progress and resolve problems of all entities 
involved in the implementation of the three TDRS mode.  This meeting will 
become a biweekly meeting until the three TDRS mode is operational.


                      FUTURE SIGNIFICANT EVENTS

The next UARS Yaw slew is scheduled for 08 August 2001 (DOY 220).  This 
maneuver will be a Forward to Reverse Flight maneuver (Southern Hemisphere 
viewing).

Attachments:  Appendix A,  Real-time Commands Sheets
              Appendix B,  Out-Of-Limits & ALERT Occurrences
              Appendix C,  Clock & Solar Array Rate Corrections
              Appendix D,  UARS Battery Performance Data
              Appendix E,  UARS Anomaly Reports

                            APPENDIX A
                      REAL-TIME COMMAND SHEETS
Week ending 06 July 2001

ORBIT     AOS          NO.       SUBSYS       REASON
-----     ---          ---        ------      ------
53603     182/1138     01-431     SOLSTICE    Power ON - Fulltime
53604     182/1415     01-432     SUSIM       Power ON - (Partial)
53607     182/1748     01-433     SUSIM       Additional Commands
53609     182/2116     01-434     ACRIM       Power ON
53618     183/1139     01-432     SUSIM       Power ON - CONTINUED
53618     183/1236     01-435     SUSIM       Power ON - COMPLETE
53619     183/1319     01-436     PEM         ZEPS / HEPS DAY-ONLY
53625     183/2339     01-437     PEM         Power ON AXIX 1 & 2
53637     184/1753     01-438     NBTR        NOOPS D-185
53640     184/2257     01-439     HALOE       Power ON
53644     185/0623     01-438     NBTR        NOOPS D-185
53644     185/0623     01-440     HALOE       Power ON complete
53646     185/0918     01-441     HALOE       SR & SS corrections
53652     185/1755     01-442     NBTR        NOOPS D-186
53658     186/0453     01-442     NBTR        NOOPS D-186
53668     186/1757     01-443     NBTR        NOOPS D-187
53673     187/0441     01-443     NBTR        NOOPS D-187
53680     187/1501     01-444     OBC/TAM     TAM UPDATE
53681     187/1611     01/443     NBTR        NOOPS D-187
53682     187/1758     01-445     NBTR        NOOPS D-188
____________________________________________________________________________

                           APPENDIX B
              OUT-OF-LIMITS & ALERTS OCCURRENCES
                  Week ending 06 July 2001

The following Out-Of-Limits occurrences were detected:

MNEMONIC     STATE     DESCRIPTION                ORBITS
--------     -----     -----------                ------
HACS4        Y-HI      GIMBLE LIMIT               53646
HAMFCAT      R-HI      TELESCOPE BARREL#3 TEMP    53646
HATBLT3      R-HI      MAIN FRAME COLD TEMP       53646


The following ALERT occurrences were detected:

MNEMONIC     DESCRIPTION     ORBITS
--------     -----------     ------
-None this report period

_______________________________________________________________________________

                           APPENDIX C
             CLOCK & SOLAR ARRAY RATE CORRECTIONS
                    Week ending 06 July 2001

                   CLOCK DRIFT RATE ADJUSTMENTS

ORBIT     TIME     CLOCK RATE CHANGE
-----     -----     -----------------
-None this report period
___________________________________________________________________________


                           APPENDIX D
                  UARS BATTERY PERFORMANCE DATA
                     Week ending 06 July 2001


     29 June 2001 - GMT Day 180 Beta = 1.7 deg, SA Pos = 269 deg

     DIFF V (mV)     TEMP (C)      CURR (amp)     EON V     AVE C/D   AVE
BAT  MAX/MIN         MAX/MIN       MAX/MIN        MIN       RATIO     DOD(%)
---  -----------     --------      ----------     ------    ------   ------
1    +728.0/+728.0   -3.05/-4.94   +0.0/+0.0      12.0      0.000     0.0
2    +11.2/-100.8    +5.47/+3.87   +28.0/-12.8    24.5      1.040     19.8
3    +22.4/-39.2     +4.51/+2.60   +27.2/-13.2    24.5      1.030     19.5


     6 July 2001 - GMT Day 187 Beta = 29.3 deg, SA Pos = 269 deg

     DIFF V (mV)     TEMP (C)      CURR (amp)     EON V     AVE C/D   AVE
BAT  MAX/MIN         MAX/MIN       MAX/MIN        MIN       RATIO     DOD(%)
---  -----------     --------      ----------     ------    -------   ------
1    +728.0/+728.0   -3.05/-4.31   +0.0/+0.0      12.0      0.000     0.0
2    +11.2/-134.4    +5.79/+4.51   +28.4/-14.4    24.8      1.030     20.7
3    +22.4/-56.0     +4.83/+3.23   +28.0/-15.2    24.8      1.020     20.7


     V/T LEVEL CHANGES
ORB#             TIME          LVL FR-TO
----             ----          ---------
-no changes      V/T 5

_______________________________________________________________________________


                                 APPENDIX E
                      UARS ANOMALY REPORTS
                    Week ending 06 July 2001

                  NEW ANOMALY REPORTS GENERATED
AR NO.     SUBSYS     ORBIT     TIME        TTR     PROBLEM/DATA LOSS
------     ------     -----     ----        ---     -----------------
01-040     HALOE       53646    185/0814    N/A     HALOE LIMITS OOL
01-041     FOT        53681     187/1634Z   14170   Generic Late Acquisition



               ANOMALY CLOSURE INFORMATION RECEIVED
AR NO.     SUBSYS     ORBIT     TIME        TTR     PROBLEM/DATA LOSS
------     ------     -----     ----        ---     -----------------
01-032     FOT        see list  June 01     14170   Generic Late Acquisition
01-040     HALOE     53646     185/0814     N/A     HALOE LIMITS OOL


             ANOMALY INVESTIGATION REPORTS GENERATED
AR NO.     SUBSYS     ORBIT     TIME        TTR     PROBLEM/DATA LOSS
------     ------     -----     ----        ---     -----------------
-None this report period


                      OPEN ANOMALY REPORTS
AR NO.     SUBSYS     ORBIT     TIME        TTR     PROBLEM/DATA LOSS
------     ------     -----     ----        ---     -----------------
01-041     FOT        see list  July 01     14170    Generic Late Acquisition


	 	SEVNTFW01/27